OpenStudy (anonymous):

At a car dealership, the manager has found that for every 18 cars sold, 5 are brought back for maintenance within the first two weeks. Next month, the manager forecasts that the dealership will sell 720 cars. What is a reasonable prediction of how many of those cars will be brought back for maintenance within the first two weeks?

OpenStudy (anonymous):

200 180 140 45

OpenStudy (anonymous):

5/18=x/720 x=(720x5)/18 x=200